OpenStudy (anonymous):

Natalie made a mistake in the following problem. Part 1) In which line was the mistake made? Part 2) Show your work and give the correct answer? Line 1 6(4) ÷ 2 − 1 + 5 Line 2 6(4) ÷ 1 + 5 Line 3 24 ÷ 1 + 5 Line 4 24 + 5 Line 5 29

OpenStudy (cwrw238):

line 2 is incorrect

OpenStudy (cwrw238):

can you see the error?

OpenStudy (anonymous):

no can you explain to me

OpenStudy (cwrw238):

6(4) should be calculated first the order of operations was wrong of you heard of PEDMAS?

OpenStudy (mayankdevnani):

every line comes to to 29 except the line one.

OpenStudy (anonymous):

what is the correct answer

OpenStudy (cwrw238):

PEDMAS is a way of remembering the order of operations Parentheses Exponential division Multiplication Addition Subtraction

OpenStudy (cwrw238):

line 2 12 - 1 + 5 = 11 + 5 = 16

OpenStudy (mayankdevnani):

Line 1 6(4) ÷ 2 − 1 + 5 =16 Line 2 6(4) ÷ 1 + 5 =29 Line 3 24 ÷ 1 + 5 =29 Line 4 24 + 5 =29 Line 5 29 =29 so the mistake was made in line 1.

OpenStudy (mayankdevnani):

see if you can follow this. Simplify parenthesis Compute powers and roots Multiply and divide Add and subtract Divide the denominator This is the right one for order of operations. as @cwrw238 says...

OpenStudy (mayankdevnani):

ok @Kathy998